please help me solve this problem. find the perimeter and area of a rectangle with a length of 6ft and a diagonal of 10 ft.
Answer by IBMath(7)

Since the length of a rectangle is 6 feet, we just need to find the width.
Since, A = lw, Perimeter = 2(l+w)
The diagonal of a rectangle splits it in half, making it into 2 right triangles. And if the hypotenuse(diagonal) is 10 ft. Just plug in the pythagorean theorem where a^2 + b^2 = c^2
So: 6^2 + b^2 = 10^2
Where is b = the width of the Rectangle and Triangle
36 + b^2 = 100
b^2 = 64
b = 8
So the area would be 8 * 6 = 48 feet
And the perimeter would be 2(8+6) = 28 feet
Good Luck!